# Superconvergent and Divergence-Free Finite Element Methods for Stokes Equation

### v1

@@ 3.1. Spaces of Lagrange multipliers
We introduce the Lagrange multiplier spaces
$\Lambda_{k}=\mathbb{P}_{k}(\mathcal{F}_{h};\mathbb{R}^{d-1})\quad\textrm{and}\quad\mathring{\Lambda}_{k}=\mathbb{P}_{k}(\mathring{\mathcal{F}}_{h};\mathbb{R}^{d-1})$
to relax the tangential-normal continuity condition in the space $\Sigma_{k}^{\rm tn}$. The embedding $E:\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\times\mathring{\Lambda}_{k}\to\mathring{M}_{k-1,k}^{-1}(\mathbb{R}^{d})$ is extended as
@@
The stress space is relaxed to $\Sigma_{k}^{-1}(\mathbb{T})=\prod_{T\in\mathcal{T}_{h}}\mathbb{P}_{k}(T;\mathbb{T})$. The weak divergence operator $\operatorname{div}_{w}$ then induces the bilinear form on $\Sigma_{k}^{-1}(\mathbb{T})\times(\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\times\mathring{\Lambda}_{k})$:
(21) $(\operatorname{div}_{w}\boldsymbol{\sigma},(\boldsymbol{v},\boldsymbol{\mu}))_{0,h}=\sum_{T\in\mathcal{T}_{h}}(\operatorname{div}\boldsymbol{\sigma},\boldsymbol{v})_{T}-\sum_{F\in\mathring{\mathcal{F}}_{h}}\!\Big(([\boldsymbol{n}^{\intercal}\boldsymbol{\sigma}\boldsymbol{n}],\,\boldsymbol{n}\!\cdot\!\boldsymbol{v})_{F}+([\Pi_{F}\boldsymbol{\sigma}\boldsymbol{n}],\,\boldsymbol{\mu})_{F}\Big),$
@@ 3.2. Weak div stability
for each $T\in\mathcal{T}_{h}$, where $\boldsymbol{n}$ is the unit outward normal on $\partial T$. Then $\Pi_{F}\boldsymbol{\tau}_{1}\boldsymbol{n}]=h_{F}\boldsymbol{\mu}_{h}$. By inverse and scaling estimates,
@@
$(\operatorname{div}_{w}\boldsymbol{\tau}_{1},(0,\boldsymbol{\mu}_{h}))_{0,h}=-\sum_{F\in\mathring{\mathcal{F}}_{h}}([\Pi_{F}\boldsymbol{\tau}_{1}\boldsymbol{n}],\boldsymbol{\mu}_{h})_{F}=\|(0,\boldsymbol{\mu}_{h})\|_{0,h}^{2},$
@@
On the other hand, by the weak divergence inf-sup (cf. (9)),
@@
The inf-sup condition (23) implies $\operatorname{div}_{w}$ is surjective. Consequently the adjoint $\operatorname{dev}{\rm grad\,}_{w}$ is injective and
$\|\operatorname{dev}{\rm grad\,}_{w}(\cdot,\cdot)\|\text{ defines a norm on }\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}\times\mathring{\Lambda}_{k}.$
Notice that the velocity space cannot be enlarged to $\mathring{\mathbb{V}}_{k,k}^{\operatorname{div}}$. But later on we will use the coercivity on the divergence free subspaces.
@@ 3.4. Error analysis
The estimate (27) demonstrates that the mixed finite element method (24) is pressure-robust.
@@ 3.5. $L^{2}$ error estimate
We refer to [29, Remark I.5.6] and [44, Section 11.5] for the $H^{2}$-regularity (29) in convex domains.
@@
Next consider case $k=0$. Applying the integration by parts to the right hand side of (31) to have
@@
Since $\operatorname{div}\tilde{\boldsymbol{u}}=0$ and $\operatorname{div}H_{0}^{2}(\Omega;\mathbb{K})=H_{0}^{1}(\Omega;\mathbb{R}^{d})\cap\ker(\operatorname{div})$ (cf. [24, Theorem 1.1]), there exists a $\tilde{\boldsymbol{\tau}}\in H_{0}^{2}(\Omega;\mathbb{K})$ such that
@@
The DoFs for space $\mathbb{V}_{h}^{d-2}$ are given in [13, (3.10)]. Let $I_{h}^{d-2}:H^{2}(\Omega;\mathbb{K})\to\mathbb{V}_{h}^{d-2}$ be the nodal interpolation operator. It holds the following commuting property:
@@
Then, employing the integration by parts and (26),
@@ 3.6. Postprocessing
We can construct a superconvergent approximation of $\boldsymbol{u}$ using the estimates (27) and (32). For each $T\in\mathcal{T}_{h}$, find $\boldsymbol{u}_{h}^{*}|_{T}\in\mathbb{P}_{k+1}(T;\mathbb{R}^{d})$ and $p_{h}^{*}\in\mathbb{P}_{k}(T)\cap L_{0}^{2}(T)$ such that
@@
(33c) $(\operatorname{div}\boldsymbol{u}_{h}^{*},q)_{T}$ $=0\qquad\qquad\qquad\;\forall\,q\in{\mathbb{P}}_{k}(T)\cap L_{0}^{2}(T),$
@@
The local problems (33) are well-posed. From (33a) we get
$Q_{0,T}(\operatorname{div}\boldsymbol{u}_{h}^{*})=Q_{0,T}(\operatorname{div}\boldsymbol{u}_{h})=0;$
combined with (33c), this yields $\operatorname{div}\boldsymbol{u}_{h}^{*}=0$ on each $T\in\mathcal{T}_{h}$. Thus $\boldsymbol{u}_{h}^{*}\in\mathbb{P}_{k+1}(\mathcal{T}_{h};\mathbb{R}^{d})$ is pointwise divergence-free elementwise.
@@ 4. Equivalent Discrete Methods
In this section, we explore several discrete methods that are equivalent to the mixed finite element scheme (24).
@@ 4.1. Hybridization of the velocity
Recall the $H^{1}$-nonconforming virtual element in [7, 15]. For integers $k\geq 0$ and $\ell=k,k-1$, the shape function space is defined as
$V_{k+1,\ell+2}^{\rm VE}(T)=\{v\in H^{1}(T):\Delta v\in\mathbb{P}_{\ell}(T),\partial_{n}v|_{F}\in\mathbb{P}_{k}(F)\textrm{ for }F\in\mathcal{F}(T)\}.$
@@ 4.2. Stabilization-free virtual element methods
For $k\geq 0$, define the space of shape functions for an $H^{1}$-nonconforming virtual element as
@@
Following the argument in [53, Section 3.1], the local virtual element space $\mathbb{V}_{k+1}^{\rm VE}(T)$ is uniquely determined by the following DoFs
@@
for all $\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{h}^{\rm VE}$ and $q_{h}\in\mathbb{P}_{k}(\mathcal{T}_{h})/\mathbb{R}$. When $k=0$, the mixed method (41) is exactly the modified Crouzeix-Raviart element method in [42, (23)].
@@
We can also use the following virtual element space for the velocity:
@@
4.3. Discrete methods based on the pseudostress-velocity-pressure formulation
@@
was employed in the MCS method [31] to discretize the traceless stress tensor $\boldsymbol{\sigma}$. However, the constraint $\Pi_{F}\boldsymbol{\tau}_{h}\boldsymbol{n}\in\mathbb{P}_{k-1}(F;\mathbb{R}^{d-1})$ prevents the discrete scheme in [31] from achieving superconvergence properties such as the estimates $\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|$ and $\|Q_{\ell}p-p_{h}\|$ in (27), and $\|I_{k,k}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}_{h}\|$ in (32).
@@ 5. Numerical examples
In this section, we present numerical examples to validate the error estimates of the mixed finite element method (24) and the post-processing procedure (33). The test problems are adapted from [30], which primarily considered higher-order cases with $k=\ell\geq 1$. Here, we focus on low-order settings with $\ell=0$ and $k=0,1$.
The computational domain is $\Omega=(0,1)^{d}$ for $d=2,3$, uniformly partitioned into simplicial meshes. All numerical experiments were developed based on the MATLAB package iiFEM [14].
@@
We first test the mixed finite element method (24) in two dimensions. Take the exact solution
$\boldsymbol{u}=\operatorname{curl}\psi_{2},\quad\text{and}\quad p=-x^{5}-y^{5}+\frac{1}{3},$
@@
Table 1. Errors $\|\boldsymbol{u}-\boldsymbol{u}_{h}\|$, $\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|$, and $\|p-p_{h}\|$ for Example 5.1 with $\ell=0$ and $k=0,1$ in two dimensions.
$h$ $(k,\ell)$ $\|\boldsymbol{u}-\boldsymbol{u}_{h}\|$ order $\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|$ order $\|p-p_{h}\|$ order
$2^{-3}$ $(0,0)$ 2.988e-03 3.103e-02 7.810e-02
$2^{-4}$ $(0,0)$ 1.284e-03 1.22 1.677e-02 0.89 3.914e-02 1.00
$2^{-5}$ $(0,0)$ 5.988e-04 1.10 8.700e-03 0.95 1.963e-02 1.00
$2^{-6}$ $(0,0)$ 2.907e-04 1.04 4.440e-03 0.97 9.840e-03 1.00
$2^{-7}$ $(0,0)$ 1.436e-04 1.02 2.247e-03 0.98 4.931e-03 1.00
$2^{-3}$ $(1,0)$ 3.296e-04 2.447e-03 7.453e-02
$2^{-4}$ $(1,0)$ 8.382e-05 1.98 6.305e-04 1.96 3.760e-02 0.99
$2^{-5}$ $(1,0)$ 2.104e-05 1.99 1.597e-04 1.98 1.880e-02 1.00
$2^{-6}$ $(1,0)$ 5.266e-06 2.00 4.016e-05 1.99 9.428e-03 1.00
$2^{-7}$ $(1,0)$ 1.317e-06 2.00 1.007e-05 2.00 4.715e-03 1.00
Table 2. Errors $\|\boldsymbol{u}-\boldsymbol{u}_{h}^{*}\|$ and $\|{\rm grad\,}_{h}(\boldsymbol{u}-\boldsymbol{u}_{h}^{*})\|$ for Example 5.1 with $\ell=0$ and $k=0,1$ in two dimensions.
$h$ $(k,\ell)$ $\|\boldsymbol{u}-\boldsymbol{u}_{h}^{*}\|$ order $\|{\rm grad\,}_{h}(\boldsymbol{u}-\boldsymbol{u}_{h}^{*})\|$ order
$2^{-3}$ $(0,0)$ 1.233e-03 2.890e-02
$2^{-4}$ $(0,0)$ 3.277e-04 1.91 1.481e-02 0.96
$2^{-5}$ $(0,0)$ 8.353e-05 1.97 7.453e-03 0.99
$2^{-6}$ $(0,0)$ 2.099e-05 1.99 3.733e-03 1.00
$2^{-7}$ $(0,0)$ 5.256e-06 2.00 1.867e-03 1.00
$2^{-3}$ $(1,0)$ 3.296e-05 2.286e-03
$2^{-4}$ $(1,0)$ 4.167e-06 2.98 5.183e-04 1.98
$2^{-5}$ $(1,0)$ 5.264e-07 2.99 1.463e-04 1.99
$2^{-6}$ $(1,0)$ 6.625e-08 2.99 3.666e-05 2.00
$2^{-7}$ $(1,0)$ 8.315e-09 2.99 9.178e-06 2.00
@@
We then test the mixed finite element method (24) in three dimensions. Take the exact solution
$\boldsymbol{u}=\operatorname{curl}(\psi_{3},\psi_{3},\psi_{3})^{\intercal},\quad\text{and}\quad p=-x^{5}-y^{5}-z^{5}+\frac{1}{2},$
@@
The numerical errors $\|\boldsymbol{u}-\boldsymbol{u}_{h}\|$, $\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|$, and $\|p-p_{h}\|$ are reported in Table 1 for Example 5.1 and Table 3 for Example 5.2, while the post-processed errors $\|\boldsymbol{u}-\boldsymbol{u}_{h}^{*}\|$ and $\|{\rm grad\,}_{h}(\boldsymbol{u}-\boldsymbol{u}_{h}^{*})\|$ are given in Table 2 for Example 5.1 and Table 4 for Example 5.2. The results in these tables demonstrate that
$\|\boldsymbol{u}-\boldsymbol{u}_{h}\|\eqsim\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|=\mathcal{O}(h^{k+1}),\quad\|p-p_{h}\|=\mathcal{O}(h),$
and for the post-processed solution $\boldsymbol{u}_{h}^{*}$,
$\|\boldsymbol{u}-\boldsymbol{u}_{h}^{*}\|=\mathcal{O}(h^{k+2}),\quad\|{\rm grad\,}_{h}(\boldsymbol{u}-\boldsymbol{u}_{h}^{*})\|=\mathcal{O}(h^{k+1}),$
which are consistent with the theoretical estimates (27) and (34)–(35).
Table 3. Errors $\|\boldsymbol{u}-\boldsymbol{u}_{h}\|$, $\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|_{0,h}$, and $\|p-p_{h}\|$ for Example 5.2 with $\ell=0$ and $k=0,1$ in three dimensions.

### v2

@@ 3.1. Spaces of Lagrange multipliers
To relax the tangential-normal continuity condition in the space $\Sigma_{k}^{\rm tn}$, we introduce the Lagrange multiplier spaces
$\Lambda_{k}=\mathbb{P}_{k}(\mathcal{F}_{h};\mathbb{R}^{d-1})\quad\textrm{and}\quad\mathring{\Lambda}_{k}=\mathbb{P}_{k}(\mathring{\mathcal{F}}_{h};\mathbb{R}^{d-1}).$
The embedding $E:\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\times\mathring{\Lambda}_{k}\to\mathring{M}_{k-1,k}^{-1}(\mathbb{R}^{d})$ is defined as
@@
The stress space is relaxed to $\Sigma_{k}^{-1}(\mathbb{T})=\mathbb{P}_{k}(\mathcal{T}_{h};\mathbb{T})$. The weak divergence operator $\operatorname{div}_{w}$ then induces the bilinear form on $\Sigma_{k}^{-1}(\mathbb{T})\times(\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\times\mathring{\Lambda}_{k})$:
(28) $(\operatorname{div}_{w}\boldsymbol{\sigma},(\boldsymbol{v},\boldsymbol{\mu}))_{0,h}=\!\sum_{T\in\mathcal{T}_{h}}(\operatorname{div}\boldsymbol{\sigma},\boldsymbol{v})_{T}-\!\sum_{F\in\mathring{\mathcal{F}}_{h}}\!\Big(([\![\boldsymbol{n}^{\intercal}\boldsymbol{\sigma}\boldsymbol{n}]\!],\,\boldsymbol{n}\cdot\boldsymbol{v})_{F}+([\![\Pi_{F}\boldsymbol{\sigma}\boldsymbol{n}]\!],\,\boldsymbol{\mu})_{F}\Big),$
@@
Again, (29) may not hold for $I_{k,k-1}^{\operatorname{div}}$, because the term $(I_{k,k}^{\operatorname{div}}\boldsymbol{v},\operatorname{div}\boldsymbol{\tau})_{T}$ requires the interior moments of the full space $\mathbb{P}_{k-1}(T;\mathbb{R}^{d})$. This property holds for RTk, but not for BDMk.
@@ 3.2. Weak div stability
for each $T\in\mathcal{T}_{h}$, where $\boldsymbol{n}$ is the unit outward normal on $\partial T$. Then $\Pi_{F}\boldsymbol{\tau}_{1}\boldsymbol{n}]\!]=h_{F}\boldsymbol{\mu}_{h}$. By inverse and scaling estimates,
@@
$(\operatorname{div}_{w}\boldsymbol{\tau}_{1},(0,\boldsymbol{\mu}_{h}))_{0,h}=-\sum_{F\in\mathring{\mathcal{F}}_{h}}([\![\Pi_{F}\boldsymbol{\tau}_{1}\boldsymbol{n}]\!],\boldsymbol{\mu}_{h})_{F}=\|(0,\boldsymbol{\mu}_{h})\|_{0,h}^{2},$
@@
On the other hand, by the weak divergence inf-sup condition (cf. (13)),
@@
The inf–sup condition (30) ensures that the weak divergence operator $\operatorname{div}_{w}$ is surjective when $\ell=k-1$. Consequently, its adjoint, the weak deviatoric gradient $\operatorname{dev}{\rm grad\,}_{w}$, is injective, and the mapping
$\|\operatorname{dev}{\rm grad\,}_{w}(\cdot,\cdot)\|$
defines a norm on the product space $\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}\times\mathring{\Lambda}_{k}$.
It is important to note that the inf-sup condition (30) may not hold for the larger space $\mathring{\mathbb{V}}_{k,k}^{\operatorname{div}}$, even though this space is important for its approximation properties. In the subsequent analysis, the coercivity of the weak operators on the divergence-free subspaces is used to resolve this conflict.
@@ 3.4. Error analysis
The estimate (34) demonstrates that the mixed finite element method (31) is pressure-robust as it depends only on the regularity of $\boldsymbol{u}$ not pp.
@@ 3.5. $L^{2}$ error estimate
We refer to [30, Remark I.5.6] and [46, Section 11.5] for the $H^{2}$-regularity (36) in convex domains.
@@
Next consider case $k=0$. Applying integration by parts to the right hand side of (38) to have
@@
Since $\operatorname{div}\tilde{\boldsymbol{u}}=0$ and $\operatorname{div}H_{0}^{2}(\Omega;\mathbb{K})=H_{0}^{1}(\Omega;\mathbb{R}^{d})\cap\ker(\operatorname{div})$ (cf. [25, Theorem 1.1]), there exists a $\tilde{\boldsymbol{\tau}}\in H_{0}^{2}(\Omega;\mathbb{K})$ such that
@@
The DoFs for space $\mathbb{V}_{h}^{d-2}$ are given in [14, (3.10)]. Let $I_{h}^{d-2}:H^{2}(\Omega;\mathbb{K})\to\mathbb{V}_{h}^{d-2}$ be the nodal interpolation operator. It holds the following commuting property:
@@
Then, employing integration by parts and (33),
@@ 3.6. Postprocessing
We can construct a superconvergent approximation of $\boldsymbol{u}$ using the estimates (34) and (39). For each $T\in\mathcal{T}_{h}$, find $\boldsymbol{u}_{h}^{*}|_{T}\in\mathbb{P}_{k+1}(T;\mathbb{R}^{d})$ and $p_{h}^{*}\in\mathbb{P}_{k}(T)/\mathbb{R}$ such that
@@
(40c) $(\operatorname{div}\boldsymbol{u}_{h}^{*},q)_{T}$ $=0\qquad\qquad\qquad\;\forall\,q\in\mathbb{P}_{k}(T)/\mathbb{R},$
@@
The local problems (40) are well-posed. From (40a), it follows that
$Q_{0,T}(\operatorname{div}\boldsymbol{u}_{h}^{*})=Q_{0,T}(\operatorname{div}\boldsymbol{u}_{h})=0.$
Combined with (40c), this identity implies that $\operatorname{div}\boldsymbol{u}_{h}^{*}=0$ for each $T\in\mathcal{T}_{h}$. Consequently, the postprocessed velocity $\boldsymbol{u}_{h}^{*}\in\mathbb{P}_{k+1}(\mathcal{T}_{h};\mathbb{R}^{d})$ is pointwise divergence-free on every element. However, we note that $\boldsymbol{u}_{h}^{*}$ is not globally $H(\operatorname{div})$-conforming, as the normal trace continuity (40a) is weakly enforced for $q\in\mathbb{P}_{0}(F)$ only.
@@
Remark 3.8.
The estimate for $\|Q_{\ell}p-p_{h}\|$ in (34) is optimal when $\ell=k$ and superconvergent when $\ell=k-1$. For fixed $k\geq 0$, since
$\mathring{\mathbb{V}}_{k,k}^{\operatorname{div}}\cap\ker(\operatorname{div})=\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}\cap\ker(\operatorname{div}),$
the discrete solutions $\boldsymbol{u}_{h}$ and $\boldsymbol{\lambda}_{h}$ are independent of $\ell$. In fact, if $(\boldsymbol{u}_{h},\boldsymbol{\lambda}_{h},p_{h})$ solves the mixed method (31) with $\ell=k$, then $(\boldsymbol{u}_{h},\boldsymbol{\lambda}_{h},Q_{k-1}p_{h})$ is the solution of the method with $\ell=k-1$. When $k\geq 1$, we have $p_{h}-Q_{k-1}p_{h}\in\mathbb{P}_{k}(T)/\mathbb{R}$ and
$\operatorname{div}\big(\mathbb{V}_{k,k}^{\operatorname{div}}(T)\cap H_{0}(\operatorname{div},T)\big)=\mathbb{P}_{k}(T)/\mathbb{R},$
so $p_{h}$ can be recovered locally from $Q_{k-1}p_{h}$. Specifically, for each $T\in\mathcal{T}_{h}$, $p_{h}-Q_{k-1}p_{h}$ is characterized by
$(\operatorname{div}\boldsymbol{v}_{h},p_{h}-Q_{k-1}p_{h})_{T}=(\boldsymbol{f},\boldsymbol{v}_{h})_{T}-(\boldsymbol{\sigma}_{h},\operatorname{dev}{\rm grad\,}_{w}(\boldsymbol{v}_{h},0))_{T}-(\operatorname{div}\boldsymbol{v}_{h},Q_{k-1}p_{h})_{T},$
for all $\boldsymbol{v}_{h}\in\mathbb{V}_{k,k}^{\operatorname{div}}(T)\cap H_{0}(\operatorname{div},T)$. We refer to [9, 54, 41] for related pressure post-processing procedures in virtual element methods.
@@ 4. Equivalent Discrete Methods
We establish connections between our mixed finite element method and several existing divergence-free schemes. We first show that hybridizing the velocity leads to an equivalent $H^{1}$-nonconforming virtual element formulation, facilitating implementation in low-order cases. Second, we demonstrate equivalence to stabilization-free virtual element methods, where the weak deviatoric gradient acts as a local projection operator. Finally, we compare our approach with the Mass-Conserving-Stress (MCS) method and show that low-order cases can be handled naturally within our framework.
@@ 4.1. Hybridization of the velocity
We recall the $H^{1}$-nonconforming virtual element in [8, 16]. For integers $k\geq 0$ and $\ell=k,k-1$, the shape function space is defined as
(44) $V_{k+1,\ell+2}^{\rm VE}(T)=\{v\in H^{1}(T):\Delta v\in\mathbb{P}_{\ell}(T),\partial_{n}v|_{F}\in\mathbb{P}_{k}(F)\textrm{ for }F\in\mathcal{F}(T)\}.$
@@ 4.2. Stabilization-free virtual element methods
For $k\geq 0$, define the space of vector-valued shape functions for an $H^{1}$-nonconforming virtual element as
@@
Following the argument in [54, Section 3.1], the local virtual element space $\mathbb{V}_{k+1}^{\rm VE}(T)$ is uniquely determined by the following DoFs
@@
for all $\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{h}^{\rm VE}$ and $q_{h}\in\mathbb{P}_{k}(\mathcal{T}_{h})/\mathbb{R}$. When $k=0$, the mixed method (49) is exactly the modified Crouzeix-Raviart element method in [44, (23)].
@@
We can also use the following virtual element space for the velocity based on (44):
@@
4.3. Mass-Conserving-Stress Method
@@
was employed in the MCS method [32] to discretize the traceless stress tensor $\boldsymbol{\sigma}$. However, the constraint $\Pi_{F}\boldsymbol{\tau}_{h}\boldsymbol{n}\in\mathbb{P}_{k-1}(F;\mathbb{R}^{d-1})$ prevents the decoupling of the error and thus not achieving superconvergence properties such as the estimates $\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|$ and $\|Q_{\ell}p-p_{h}\|$ in (34), and $\|I_{k,k}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}_{h}\|$ in (39).
@@ 5. Numerical Examples
In this section, we verify the error estimates for the mixed finite element method (31) and the postprocessing scheme (40). Both smooth and singular test problems are considered, with focus on the low-order cases $\ell=0$ and $k=0,1$. The smooth benchmark problems are adapted from [31], where examples for $k=\ell\geq 1$ are presented. Throughout, the computational domains are discretized by uniform simplicial meshes, and all computations are carried out in MATLAB using iiFEM [15].
5.1. Smooth problems with non-slip boundary conditions
@@
We first test the mixed finite element method (31) in two dimensions on the unit square domain $\Omega=(0,1)^{2}$. The exact solution is
$\boldsymbol{u}=\operatorname{curl}\psi_{2},\qquad p=-x^{5}-y^{5}+\frac{1}{3},$
@@
Table 2. Errors for Example 5.1 with $\ell=0$ and $k=0,1$ in two dimensions.
$h$ $(k,\ell)$ $\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|_{0,h}$ order $\|p-p_{h}\|$ order $\|\boldsymbol{u}-\boldsymbol{u}_{h}^{*}\|$ order $\|\nabla_{h}(\boldsymbol{u}-\boldsymbol{u}_{h}^{*})\|$ order
$2^{-3}$ $(0,0)$ 3.103e-02 7.810e-02 1.233e-03 2.890e-02
$2^{-4}$ $(0,0)$ 1.677e-02 0.89 3.914e-02 1.00 3.277e-04 1.91 1.481e-02 0.96
$2^{-5}$ $(0,0)$ 8.700e-03 0.95 1.963e-02 1.00 8.353e-05 1.97 7.453e-03 0.99
$2^{-6}$ $(0,0)$ 4.440e-03 0.97 9.840e-03 1.00 2.099e-05 1.99 3.733e-03 1.00
$2^{-7}$ $(0,0)$ 2.247e-03 0.98 4.931e-03 1.00 5.256e-06 2.00 1.867e-03 1.00
$2^{-3}$ $(1,0)$ 2.447e-03 7.453e-02 3.296e-05 2.286e-03
$2^{-4}$ $(1,0)$ 6.305e-04 1.96 3.760e-02 0.99 4.167e-06 2.98 5.183e-04 1.98
$2^{-5}$ $(1,0)$ 1.597e-04 1.98 1.880e-02 1.00 5.264e-07 2.99 1.463e-04 1.99
$2^{-6}$ $(1,0)$ 4.016e-05 1.99 9.428e-03 1.00 6.625e-08 2.99 3.666e-05 2.00
$2^{-7}$ $(1,0)$ 1.007e-05 2.00 4.715e-03 1.00 8.315e-09 2.99 9.178e-06 2.00
@@
We next consider the mixed finite element method (31) in three dimensions on the unit cube domain $\Omega=(0,1)^{3}$. The exact solution is given by
$\boldsymbol{u}=\operatorname{curl}(\psi_{3},\psi_{3},\psi_{3})^{\intercal},\qquad p=-x^{5}-y^{5}-z^{5}+\frac{1}{2},$
@@
Table 3. Errors for Example 5.2 with $\ell=0$ and $k=0,1$ in three dimensions.
$h$ $(k,\ell)$ $\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|_{0,h}$ order $\|p-p_{h}\|$ order $\|\boldsymbol{u}-\boldsymbol{u}_{h}^{*}\|$ order $\|\nabla_{h}(\boldsymbol{u}-\boldsymbol{u}_{h}^{*})\|$ order
$2^{-1}$ $(0,0)$ 9.787e-03 2.942e-01 4.167e-04 4.576e-03
$2^{-2}$ $(0,0)$ 4.960e-03 1.06 1.649e-01 0.84 1.565e-04 1.41 2.880e-03 0.67
$2^{-3}$ $(0,0)$ 2.431e-03 0.95 8.501e-02 0.96 4.400e-05 1.83 1.539e-03 0.90
$2^{-4}$ $(0,0)$ 1.260e-03 0.95 4.284e-02 0.99 1.143e-05 1.95 7.836e-04 0.97
$2^{-5}$ $(0,0)$ 6.397e-04 0.98 2.146e-02 1.00 2.889e-06 1.98 3.937e-04 0.99
$2^{-1}$ $(1,0)$ 2.431e-03 2.942e-01 9.624e-05 1.894e-03
$2^{-2}$ $(1,0)$ 5.715e-04 2.09 1.649e-01 0.84 1.493e-05 2.69 5.478e-04 1.79
$2^{-3}$ $(1,0)$ 1.541e-04 1.89 8.501e-02 0.96 2.035e-06 2.88 1.486e-04 1.88
$2^{-4}$ $(1,0)$ 3.975e-05 1.96 4.284e-02 0.99 2.627e-07 2.95 3.808e-05 1.96
Numerical errors for the smooth problems (Examples 5.1 and 5.2) are reported in Tables 2 and 3. The observed rates agree with the theoretical estimates (34) and (41)–(42). In particular,
$\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|_{0,h}=\mathcal{O}(h^{k+1}),\qquad\|p-p_{h}\|=\mathcal{O}(h),$
and the postprocessed velocity satisfies
$\|\boldsymbol{u}-\boldsymbol{u}_{h}^{*}\|=\mathcal{O}(h^{k+2}),\qquad\|\nabla_{h}(\boldsymbol{u}-\boldsymbol{u}_{h}^{*})\|=\mathcal{O}(h^{k+1}).$
5.2. Singular problem with slip boundary conditions
We next consider a singular test problem on the two-dimensional L-shaped domain $\Omega=(-1,1)^{2}\setminus\big([0,1]\times[-1,0]\big)$ with boundary condition
$\boldsymbol{u}\cdot\boldsymbol{n}=0\quad\text{and}\quad\boldsymbol{t}^{\intercal}\boldsymbol{\sigma}\boldsymbol{n}=0\quad\text{on }\partial\Omega.$
The re-entrant corner at the origin induces reduced regularity of the exact solution and hence reduced convergence rates on uniform meshes.
We impose this boundary condition by using the space $\mathring{\Sigma}_{k}^{\rm tn}$. See Remark 2.7.
Example 5.3.
Let $(r,\theta)$ denote polar coordinates centered at the re-entrant corner of $\Omega$. The corner is located at $(0,0)$, and $\theta\in(0,3\pi/2)$ in $\Omega$. Define,
$\Phi(x,y)=(1-x^{2})^{3}(1-y^{2})^{3},\qquad\Psi_{s}(r,\theta)=r^{1+\lambda}\sin((\lambda-1)\theta),\qquad\lambda=\tfrac{1}{3},$
and set
$\boldsymbol{u}={\rm curl\,}(\Phi\Psi_{s}).$
Further, let
$g(\theta)=\sin((\lambda-1)\theta),\quad Q(\theta)=(1+\lambda)^{2}g(\theta)+g^{\prime\prime}(\theta),\quad p_{s}(r,\theta)=\frac{1}{\lambda-1}r^{\lambda-1}Q^{\prime}(\theta).$
The pressure is defined by
$p(x,y)=\Phi(x,y)p_{s}(r,\theta)-c_{0},\quad c_{0}=|\Omega|^{-1}\int_{\Omega}\Phi p_{s}\,dx,$
so that $p\in L_{0}^{2}(\Omega)$.
(a) $\boldsymbol{u}_{h}\text{ for }k=0,\ell=0$
(b) $\boldsymbol{u}_{h}\text{ for }k=1,\ell=0$
(c) Exact solution $\boldsymbol{u}$
Figure 2. Flow structure of the velocity field $\boldsymbol{u}$ on the L-shaped domain. Left and middle: numerical solutions. Right: exact solution.
Table 4. Errors for Example 5.3 with $\ell=0$ and $k=0,1$ in the L-shaped domain.